The Programme for International Student Assessment found that 48% of elementary school aged girls reported that they were not confident in their abilities in math. Jay is interested in finding out whether this proportion is higher for women in college. He conducts a study where the conditions of inference are met and obtains the following results, using an α=0.05α=0.05 significance level.

Question 2: What should Jay's alternative hypothesis be?

Select one:

a. p^>0.48p^>0.48

b. p^>0.545p^>0.545

c. p>0.48p>0.48

d. p>0.545p>0.545

e. p=0.48p=0.48

f. p^≠0.48p^≠0.48

g. p≠0.545p≠0.545

Question 3: What is the appropriate standard error to use when conducting the above test?

Select one:

a. 0.0353

b. 0.001

c. 0.0339

d. 0.048

Question 4

Not yet answered

Points out of 3.00

Flag question

Question text

Which of the following is an accurate interpretation of the P-value Jay found?

Select one:

a. It is not significant, so it has no interpretation.

b. It is the probability that the true population proportion is 0.48.

c. It is the probability of finding a sample proportion of 0.545 or larger, given that the true population proportion is 0.48.

d. It is the true proportion of college women who report that they are confident in their math abilities.
Answer #1

2)c. p>0.48

3)

standard error =sqrt(p(1-p)/n)=sqrt(0.48*0.52/200) =0.0353

4)

c. It is the probability of finding a sample proportion of 0.545 or larger, given that the true population proportion is 0.48.

(since p value is probabiltiy of getting as or more extreme test statistic given null is true)